What us the equation of the circle touching x=0,y=0 andx=2

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
PUTTING C=2 YOU CAN GET YOUR ANSWER.
The circle touches X axis, Y axis and a line 
parallel to Y axis at a distance c from the y axis. 
∴ its radius is = c/2 and the centre 
has to be either (c/2, c/2 ) or (-c/2, -c/2) 
The equation is (x + c/2)² + (y + c/2)² = (c/2)² 
and /or (x - c/2)² + (y - c/2)² = (c/2)² 
These are two circles that will meet the above conditions.